OpenStudy (anonymous):

1/5 % when converted to fraction is _____ ??

OpenStudy (dumbcow):

when converting percent to fraction divide by 100 \[\frac{\frac{1}{5}}{100}\] \[=\frac{1}{5}*\frac{1}{100} = \frac{1}{500}\]